OpenStudy (anonymous):

Calculate the Maturity Value of a loan if borrowing $6,000 at 7% for 48 months.

OpenStudy (anonymous):

I=p*r*t where p=principal, r=rate, and t=time. I think that's the formula XD

OpenStudy (anonymous):

I=interest

OpenStudy (anonymous):

so what should i do? :o

OpenStudy (anonymous):

I think you go like I=$6,000*7%*48

OpenStudy (anonymous):

So I think I=$20,160.

OpenStudy (wolf1728):

If you borrow 6,000 for 48 months at 8% then the monthly payment will be 143.68. Is the maturity value 48*143.68? or 6,896.64

OpenStudy (wolf1728):

Typo - I meant 7%

OpenStudy (anonymous):

They gave me this options: a. 7,680 b. 8,100 c. 1,680 d. 6,840

OpenStudy (wolf1728):

Well 6,840 is close to what I calculated - but it would be nice if it came out exactly.

OpenStudy (anonymous):

yeah :/

OpenStudy (wolf1728):

I'm looking up the "maturity value" definition right now.

OpenStudy (wolf1728):

I found a formula here: http://www.wikihow.com/Calculate-Maturity-Value Maturity Value = Principal * (1+monthly rate)^# compounding periods 7% yearly (or .07) = 0.0058333333 monthly Maturity Value = $6,000 * (1.0058333333)^48 Maturity Value = $6,000 * 1.3220538779 Maturity Value = 7,932.32 and that doesn't come close to the 4 choices either